OpenStudy (anonymous):

Solve for x. 3:12 = x:16 2 1/4 4 64 Can anyone help with this? I don't understand how to do it V___V

OpenStudy (fifciol):

\[\frac{ 3 }{12 }=\frac{x }{ 16 }\] \[3*16=x*12\] x=4

OpenStudy (ddcamp):